OpenStudy (anonymous):

Despite an historical "no layoff policy, FedEx laid off 900 employees in the early part of 2009 as reported the February 9, 2009, Memphis Business Journal. If FedEx (FDX) head quartered in Memphis Tennessee has 233,457 U.S. employees what percent were laid off in February? Round to the nearest tenth percent. We have to use the formula new value - Previous value /Previous value x 100

OpenStudy (anonymous):

I took the 233,457 - 900 = 232557 /900 x 100 = I got 25839.67% I believe I might have done something wrong here but you can't take 900 - 233,457